### v2

# Optimal Control of Nonlinear Systems with Unknown Dynamics

Abstract
This paper presents a data-driven method for finding a closed-loop optimal controller, which minimizes a specified infinite-horizon cost function for systems with unknown dynamics given any arbitrary initial state. Suppose the closed-loop optimal controller can be parameterized by a given class of functions, hereafter referred to as the policy. The proposed method introduces a novel gradient estimation framework, which approximates the gradient of the cost function with respect to the policy parameters via integrating the Koopman operator with the classical concept of actor-critic. This enables the policy parameters to be tuned iteratively using gradient descent to achieve an optimal controller, leveraging the linearity of the Koopman operator. The convergence analysis of the proposed framework is provided. The effectiveness of the method is demonstrated through comparisons with a model-free reinforcement learning approach, and its control performance is further evaluated through simulations against model-based optimal control methods that solve the same optimal control problem utilizing the exact system dynamics.
{IEEEkeywords}
Actor-Critic Algorithm, Koopman Operator, Optimal Control, Unknown Dynamics.
1 Introduction
[…]
Optimal control theory provides a foundational mathematical framework for the design of control strategies aimed at minimizing user-defined cost functions, typically under the assumption of known system dynamics [1]. For systems with unknown dynamics, model-free methods have emerged as a promising alternative for deriving the closed-loop optimal controller without using the system dynamics [2]. These approaches commonly assume unknown dynamics as a Markov decision process and use parameterized policies as closed-loop controllers. They are often implemented using an actor-critic structure [3], where the critic evaluates policy performance using observed data, and the actor updates the policy parameters based on the critic’s feedback. Here, the critic approximates the expectation of the cost function by minimizing the temporal difference error [4] without using the system dynamics, while the actor optimizes the policy by minimizing the critic, which is also referred to as policy gradient methods [5]. Notable contributions in this domain include the deep Q network [6], and proximal policy optimization [7], which are well-suited for discrete action spaces. For continuous action spaces, deterministic policy gradient algorithms [8] have been incorporated into the deep Q network framework [9]. Nevertheless, actor-critic methods typically require extensive trial-and-error iterations to identify the closed-loop optimal controllers [10], which can present challenges in real-world applications.
To improve the efficiency of model-free methods in identifying closed-loop optimal controllers, recent attention has been given to data-driven model-based methods, which involve estimating system dynamics and subsequently deriving closed-loop optimal controllers using the learned models. For example, integral reinforcement learning [11] directly determines optimal feedback gains from input-output data collected along system trajectories, typically assuming linear system dynamics. For systems with completely unknown nonlinear dynamics, several studies [12, 13, 14, 15, 16] focus on learning unknown dynamics using neural networks, followed by leveraging classical model-based control strategies such as model predictive control (MPC) to compute optimal control actions. In the area of model-based reinforcement learning [17], popular methods [18, 19] leverage Gaussian processes to approximate system dynamics, while works [20, 21] utilize deep neural networks (DNNs) to learn the dynamics. These techniques facilitate the generation of closed-loop optimal controllers by performing policy gradient updates using the learned models to directly minimize the user-defined cost function. Despite the successful integration of dynamic learning and policy design, several challenges persist, particularly due to the nonlinearity of learned dynamic models. These challenges include difficulties in verifying critical properties of the learned dynamics, such as controllability and stability, as noted in [22]. Furthermore, the significant nonlinearity in the learned dynamics often results in substantial computational complexity, which complicates practical implementation [23].
The Koopman operator [24] offers an alternative approach to approximate nonlinear systems with linear dynamics based on state-control pairs [25, 26, 27], enabling the evaluation of key properties of the learned dynamics, such as observability and controllability. Techniques like extended dynamic mode decomposition [28] transform the state space into a higher-dimensional space, making the dynamics approximately linear through carefully selected lifting functions. To address the complexity of choosing these functions, recent studies, for instance, [29], has employed deep learning techniques to discover the eigenfunctions of the Koopman operator. Deep Koopman operator (DKO) methods [30, 31, 32, 33] utilize DNNs as lifting functions, optimizing them with state-control pairs and a properly defined loss function. Model-based controllers, such as MPC, can then be integrated with the learned Koopman dynamics model [30]. Nevertheless, inaccuracies in the learned model may result in cumulative errors during cost function computation when propagating system states, thereby affecting the reliability of the approach.
Motivated by the above, this paper proposes a data-driven framework to derive the closed-loop optimal controller leveraging the combined benefits of actor-critic algorithms and dynamics learning using the DKO approach. This is achieved by decomposing the bilevel optimization problem of the actor-critic structure through the integration of the DKO into the framework, enabling independent tuning of the dynamics approximator and the critic. The primary contribution of this paper lies in developing a parameter-tuning framework that simultaneously approximates unknown dynamics, refines the critic, and optimizes the parameterized policy based on the critic’s evaluation using a one-time-step prediction from the DKO. The proposed approach is distinct from recent advancements in integrating the Koopman operator into reinforcement learning (RL). For instance, [34] employs the Koopman operator to augment the static offline datasets during training, whereas the present work focuses on estimating the policy gradient using the Koopman operator. Similarly, [35] leverages the Koopman operator to directly approximate the critic, enhancing the actor-critic framework, while the proposed method primarily leverages the DKO to predict future system states, facilitating improved policy optimization.
The remainder of the paper is organized as follows. Section 2 formulates the problem. Section 3 introduces the proposed framework and offers the corresponding theoretical analysis. Section 4 details an online algorithm for efficient framework implementation and validates the framework and theoretical findings through numerical simulations. Finally, Section 5 provides concluding remarks.
[…]
2 Problem
Consider the following discrete-time dynamical system:
$\boldsymbol{x}(t+1)=\boldsymbol{f}(\boldsymbol{x}(t),\boldsymbol{u}(t)),\quad\boldsymbol{x}(0)=\boldsymbol{x}_{0},$ (1)
where $t=0,1,2,\cdots$ denotes the time index, $\boldsymbol{x}(t)\in\mathcal{X}\subset\mathbb{R}^{n}$ and $\boldsymbol{u}(t)\in\mathcal{U}\subset\mathbb{R}^{m}$ denote the system state and control input at time $t$, respectively, $\boldsymbol{x}_{0}$ is the given arbitrary vector, and $\boldsymbol{f}:\mathcal{X}\times\mathcal{U}\rightarrow\mathcal{X}$ denotes the time-invariant, Lipschitz continuous, and unknown dynamics mapping. We assume $\mathcal{X}$ is a countably infinite state space. Notably, for the remainder of this paper, we distinguish between system state-input variables and fixed observed state-input pairs by denoting $\boldsymbol{x}_{t}$ and $\boldsymbol{u}_{t}$ as constant state and input vectors observed at time $t$, respectively.
At any time $t^{*}\in\{0,1,2,\cdots\}$, let $\boldsymbol{x}(t^{*})=\boldsymbol{x}_{t^{*}}$, an infinite-horizon cost function under the unknown dynamics in (1) is defined as follows:
$J_{\mu,t^{*}}=\sum_{t=t^{*}}^{\infty}\gamma^{t-t^{*}}c(\boldsymbol{x}(t),\boldsymbol{u}(t)),$ (2)
where $c:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}$ denotes a bounded and Lipschitz continuous stage cost function, $10<\gamma<$ 1 is a discount factor, and the control inputs are assumed to be chosen from a policy obeying the following form:
$\boldsymbol{u}(t)=\boldsymbol{\mu}(\boldsymbol{x}(t),\boldsymbol{\theta}^{\mu}),$ (3)
where $\boldsymbol{\mu}(\cdot,\boldsymbol{\theta}^{\mu}):\mathcal{X}\rightarrow\mathcal{U}$ denotes a known Lipschitz continuous function with a tunable parameter $\boldsymbol{\theta}^{\mu}\in\mathbb{R}^{q}$.
Assume that the closed-loop optimal controller that minimizes $J_{\mu,t^{*}}$ in (2) can be represented by $\boldsymbol{\mu}(\boldsymbol{x}(t),\boldsymbol{\theta}^{\mu*})$. The problem of interest is to find a data-driven framework to achieve $\boldsymbol{\theta}^{\mu*}$. Correspondingly, given any arbitrary $\boldsymbol{x}(t^{*})=\boldsymbol{x}_{t^{*}}$, this paper aims to solve the following optimization problem without knowing the actual dynamics $\boldsymbol{f}$ in (1):
[…]
Remark 1
The optimization problem in (4) is a well-established optimal control problem when the dynamics $\boldsymbol{f}$ in (1) is known [36]. In the case where $\boldsymbol{f}$ is unknown, one approach to solve (4) involves model-free methods, such as reinforcement learning techniques, which typically require a substantial amount of data to achieve an optimal solution. In the following section, we will propose a framework that concurrently approximates $\boldsymbol{f}$ using the Koopman operator and optimizes $\boldsymbol{\theta}^{\mu}$ to find the $\boldsymbol{\theta}^{\mu*}$. The method is shown to have better efficiency of computing $\boldsymbol{\theta}^{\mu*}$ in later simulations.
3 Main Results
In this section, we first identify the primary challenges and fundamental concepts underlying the proposed framework. Afterward, we introduce a data-driven tuning framework to solve (4). Finally, we present the convergence analysis of the proposed framework.
3.1 Challenges and Key Ideas
An iterative method to find $\boldsymbol{\theta}^{\mu*}$ in (4) is to employ the gradient descent method. Specifically, let $k=0,1,2,\cdots$ denote the iteration index, $\boldsymbol{\theta}_{k}^{\mu}\in\mathbb{R}^{q}$ denote the estimation of $\boldsymbol{\theta}^{\mu*}$ at the $k$-th iteration, and $\alpha_{k}^{\mu}$ represent the step size at the $k$-th iteration corresponding to $\boldsymbol{\theta}_{k}^{\mu}$. Given an arbitrary initial $\boldsymbol{\theta}_{0}^{\mu}$, the parameter $\boldsymbol{\theta}_{k}^{\mu}$ is updated iteratively using the gradient of $J_{\mu,t^{*}}$ as follows:
$\boldsymbol{\theta}_{k+1}^{\mu}=\boldsymbol{\theta}_{k}^{\mu}-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}(\boldsymbol{\theta}_{k}^{\mu}),\quad\boldsymbol{\theta}_{0}^{\mu}\ \text{given},$ (5)
where
$\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}=$ $\nabla_{\boldsymbol{\theta}^{\mu}}c(\boldsymbol{x}(t^{*}),\boldsymbol{u}(t^{*}))+\gamma\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}+1},$ (6)
$=$ $\frac{\partial c_{t^{*}}}{\partial\boldsymbol{u}_{t^{*}}}\frac{\partial\boldsymbol{u}_{t^{*}}}{\partial\boldsymbol{\theta}^{\mu}}+\sum_{t=t^{*}+1}^{\infty}\gamma^{t-t^{*}}(\frac{\partial c_{t}}{\partial\boldsymbol{u}(t)}\frac{\partial\boldsymbol{u}(t)}{\partial\boldsymbol{\theta}^{\mu}}$
$+\frac{\partial c_{t}}{\partial\boldsymbol{x}(t)}\frac{\partial\boldsymbol{x}(t)}{\partial\boldsymbol{u}(t-1)}\frac{\partial\boldsymbol{u}(t-1)}{\partial\boldsymbol{\theta}^{\mu}}).$
Here, $c_{t}\coloneqq c(\boldsymbol{x}(t),\boldsymbol{u}(t))$ is introduced for notation brevity, and the selection of $\alpha_{k}^{\mu}$ will be discussed in detail later in this paper. Throughout this paper, we refer to $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$ as the policy gradient.
Two key challenges arise when computing the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$ in (6) due to the unknown dynamics $f$ in (1). First, the system dynamics $\boldsymbol{f}$ is assumed to be unknown in the present work, making the gradient $\frac{\partial\boldsymbol{x}(t)}{\partial\boldsymbol{u}(t-1)}$ unknown as well. Second, calculating $\sum_{t=t^{*}+1}^{\infty}\frac{\partial c_{t}}{\partial\boldsymbol{x}(t)}\frac{\partial\boldsymbol{x}(t)}{\partial\boldsymbol{u}(t-1)}\frac{\partial\boldsymbol{u}(t-1)}{\partial\boldsymbol{\theta}^{\mu}}+\frac{\partial c_{t}}{\partial\boldsymbol{u}(t)}\frac{\partial\boldsymbol{u}(t)}{\partial\boldsymbol{\theta}^{\mu}}$ requires future system states $\boldsymbol{x}(t^{*}+1),\boldsymbol{x}(t^{*}+2),\cdots$ that evolve according to system dynamics. Even if one can approximate the dynamics $\boldsymbol{f}$ with small estimation errors, these errors may still accumulate over an infinite time horizon, potentially resulting in inaccuracies in calculating $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$.
To address the two challenges in computing the policy gradient, we propose a framework that applies the deep Koopman operator (DKO) to approximate the unknown dynamics $\boldsymbol{f}$ in a linear form, enabling efficient approximation of $\frac{\partial\boldsymbol{x}(t+1)}{\partial\boldsymbol{u}(t)}$ and potentially improving gradient-based optimization. Subsequently, we use the temporal difference error technique [37] to approximate $J_{\mu,t^{*}+1}$ such that $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}+1}$ can be estimated using only a one-time-step prediction from the DKO to reduce the computational complexity.
3.2 The Proposed Framework
We now introduce a data-driven framework to approximate the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$ in (6) such that one can solve (4) by tuning $\boldsymbol{\theta}^{\mu}$ following (5). The proposed gradient estimation framework comprises three components, updated concurrently: (i) a DKO block for approximating the unknown system dynamics and enabling future state predictions; (ii) a critic block for estimating the cost function $J_{\mu,t^{*}}$ in (2); and (iii) an actor block for optimizing the policy based on the critic’s evaluation, utilizing a one-time-step prediction from the DKO.
To proceed, we consider a given dataset consisting of all observed tuples, represented as $\mathcal{D}=\cup_{i=1}^{N}\{(\boldsymbol{x}_{i},\boldsymbol{u}_{i},\boldsymbol{x}_{i}^{+})\}$, with its index set denoted by $\mathcal{I}_{D}=\{1,2,\cdots,N\}$. Here, $\boldsymbol{x}_{i}\coloneqq\boldsymbol{x}_{t_{i}}$ and $\boldsymbol{u}_{i}\coloneqq\boldsymbol{u}_{t_{i}}$ represent the system state and control input observed at time $t_{i}$, respectively, while $\boldsymbol{x}_{i}^{+}\coloneqq\boldsymbol{x}_{t_{i}+1}$ denotes the resulting system state obtained from $\boldsymbol{x}_{i}$ after applying $\boldsymbol{u}_{i}$ to the unknown dynamics in (1). This notation is introduced because the proposed framework does not require the tuples to be sequentially connected, meaning $t_{i}+1$ is not necessarily equal to $t_{i+1}$. Furthermore, we assume that $\boldsymbol{u}_{i}$ is sampled from some probability distribution with non-zero probability over all actions, which we refer to as the behavior policy.
Dynamics approximation using deep Koopman operator (DKO). To approximate the unknown dynamics $\boldsymbol{f}$ in (1), this paper utilizes the DKO to obtain the following estimated dynamics:
$\boldsymbol{x}(t+1)=C^{*}\Big{(}A^{*}\boldsymbol{g}(\boldsymbol{x}(t),\boldsymbol{\theta}^{f*})+B^{*}\boldsymbol{u}(t)\Big{)},$ (7)
where $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f*}):\mathcal{X}\rightarrow\mathbb{R}^{r},r\geq n$ is a Lipschitz continuous function typically represented by a DNN parameterized by $\boldsymbol{\theta}^{f*}\in\mathbb{R}^{p}$, and $A^{*}\in\mathbb{R}^{r\times r},B^{*}\in\mathbb{R}^{r\times m},C^{*}\in\mathbb{R}^{n\times r}$ are constant matrices. Here, the estimated dynamics (7) is derived based on the following DKO representation:
$\boldsymbol{g}(\boldsymbol{x}(t+1),\boldsymbol{\theta}^{f*})$ $=A^{*}\boldsymbol{g}(\boldsymbol{x}(t),\boldsymbol{\theta}^{f*})+B^{*}\boldsymbol{u}(t),$ (8)
$\boldsymbol{x}(t+1)$ $=C^{*}\boldsymbol{g}(\boldsymbol{x}(t+1),\boldsymbol{\theta}^{f*}),$ (9)
where (8) describes the evolution of the system dynamics in the lifted space and (9) further assumes the existence of a linear mapping between $\boldsymbol{x}(t+1)$ and its lifted states $\boldsymbol{g}(\boldsymbol{x}(t+1),\boldsymbol{\theta}^{f*})$. We refer to existing work [33] for a detailed analysis of the estimation errors of (7).
To achieve the estimated dynamics in (7), we define the following optimization problem based on the dataset $\mathcal{D}$:
$A^{*},B^{*},C^{*},\boldsymbol{\theta}^{f*}=\arg\min_{A,B,C,\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A,B,C,\boldsymbol{\theta}^{f}),$
where
${\mathbf{L}}_{f}=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}(\underbrace{\frac{1}{2}\parallel\boldsymbol{g}(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{f})-A\boldsymbol{g}(\boldsymbol{x}_{i},\boldsymbol{\theta}^{f})-B\boldsymbol{u}_{i}\parallel^{2}}_{\delta_{i}(A,B,\boldsymbol{\theta}^{f})}$ (10)
$+\underbrace{\frac{1}{2}\parallel\boldsymbol{x}_{i}^{+}-C\boldsymbol{g}(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{f})\parallel^{2}}_{\bar{\delta}_{i}(C,\boldsymbol{\theta}^{f})}).$
Here, $\delta_{i}(A,B,\boldsymbol{\theta}^{f})$ and $\bar{\delta}_{i}(C,\boldsymbol{\theta}^{f})$ are designed to approximate (8) and (9), respectively. Observing (10) involves four variables, a practical way to minimize ${\mathbf{L}}_{f}$ is first to establish the relationship between $A$, $B$, $C$ and $\boldsymbol{\theta}^{f}$. This enables the computation of the gradient of ${\mathbf{L}}_{f}$ with respect to $\boldsymbol{\theta}^{f}$, thereby allowing the application of gradient descent for optimization.
To this end, let $\boldsymbol{\theta}_{k}^{f}$ be the estimation of $\boldsymbol{\theta}^{f*}$ at the $k$-th iteration, we introduce the following data matrices constructed using $\mathcal{D}$:
${\mathbf{X}}$ $=[\boldsymbol{x}_{1},\boldsymbol{x}_{2},\cdots,\boldsymbol{x}_{N}]\in\mathbb{R}^{n\times N},$ (11)
$\mathbf{U}$ $=[\boldsymbol{u}_{1},\boldsymbol{u}_{2},\cdots,\boldsymbol{u}_{N}]\in\mathbb{R}^{m\times N},$
$\bar{{\mathbf{X}}}$ $=[\boldsymbol{x}_{1}^{+},\boldsymbol{x}_{2}^{+},\cdots,\boldsymbol{x}_{N}^{+}]\in\mathbb{R}^{n\times N},$
${\mathbf{G}}_{k}$ $=[\boldsymbol{g}(\boldsymbol{x}_{1},\boldsymbol{\theta}_{k}^{f}),\boldsymbol{g}(\boldsymbol{x}_{2},\boldsymbol{\theta}_{k}^{f}),\cdots,\boldsymbol{g}(\boldsymbol{x}_{N},\boldsymbol{\theta}_{k}^{f})]\in\mathbb{R}^{r\times N},$
$\bar{{\mathbf{G}}}_{k}$ $=[\boldsymbol{g}(\boldsymbol{x}_{1}^{+},\boldsymbol{\theta}_{k}^{f}),\boldsymbol{g}(\boldsymbol{x}_{2}^{+},\boldsymbol{\theta}_{k}^{f}),\cdots,\boldsymbol{g}(\boldsymbol{x}_{N}^{+},\boldsymbol{\theta}_{k}^{f})]\in\mathbb{R}^{r\times N}.$
If the matrices $\bar{{\mathbf{G}}}_{k}\in\mathbb{R}^{r\times N}$ and $\begin{bmatrix}{\mathbf{G}}_{k}\\ \mathbf{U}\end{bmatrix}\in\mathbb{R}^{(r+m)\times N}$ in (11) are with full row rank, meaning they are right-invertible, we propose the following updating rule for $\boldsymbol{\theta}_{k}^{f}$:
$\boldsymbol{\theta}_{k+1}^{f}=\boldsymbol{\theta}_{k}^{f}-\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f}),\quad\boldsymbol{\theta}_{0}^{f}\ \text{given},$ (12)
where
$=\arg\min_{[A\ B]}\sum_{i\in\mathcal{I}_{D}}\delta_{i}(A,B,\boldsymbol{\theta}_{k}^{f})=\bar{{\mathbf{G}}}_{k}\begin{bmatrix}{\mathbf{G}}_{k}\\ \mathbf{U}\end{bmatrix}^{\dagger},$ (13)
$C_{k}$ $=\arg\min_{C}\sum_{i\in\mathcal{I}_{D}}\bar{\delta}_{i}(C,\boldsymbol{\theta}_{k}^{f})=\bar{{\mathbf{X}}}\bar{{\mathbf{G}}}_{k}^{\dagger},$
are constant matrices determined by $\boldsymbol{\theta}_{k}^{f}$, and
$\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}(\nabla_{\boldsymbol{\theta}^{f}}\delta_{i}(A_{k},B_{k},\boldsymbol{\theta}_{k}^{f})+\nabla_{\boldsymbol{\theta}^{f}}\bar{\delta}_{i}(C_{k},\boldsymbol{\theta}_{k}^{f})).$
In the subsequent subsection (Section 3.3), we present the convergence analysis of the update rule in (12) and explain why the matrices updating rule in (13) offers advantages over using arbitrarily assigned constant matrices.
Cost function approximation (critic). A common approach for approximating $J_{\mu,t^{*}}$ in (2) under a fixed control policy is to find a parameterized function $V(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{J*})=J_{\mu,t^{*}}$, where $V(\cdot,\boldsymbol{\theta}^{J}):\mathcal{X}\rightarrow\mathbb{R}$ is a Lipschitz continuous function with a known structure and tunable parameter $\boldsymbol{\theta}^{J}\in\mathbb{R}^{s}$. Given $\mathcal{D}$, the parameter $\boldsymbol{\theta}^{J*}$ can be obtained by minimizing the following loss function, formulated based on the temporal difference (TD) learning [37]:
${\mathbf{L}}_{J}(\boldsymbol{\theta}^{J})=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}\underbrace{\frac{1}{2}(c(\boldsymbol{x}_{i},\boldsymbol{u}_{i})+\gamma V(\boldsymbol{x}_{i}^{+},\boldsymbol{\theta}^{J})\!-\!V(\boldsymbol{x}_{i},\boldsymbol{\theta}^{J}))^{2}}_{\delta_{i}^{TD}(\boldsymbol{\theta}^{J})}$
using the gradient descent method
$\boldsymbol{\theta}_{k+1}^{J}=\boldsymbol{\theta}_{k}^{J}-\alpha_{k}^{J}\nabla_{\boldsymbol{\theta}^{J}}{\mathbf{L}}_{J}(\boldsymbol{\theta}_{k}^{J}),\quad\boldsymbol{\theta}_{0}^{J}\ \text{given},$ (14)
where $\boldsymbol{\theta}_{k}^{J}$ is the estimation of $\boldsymbol{\theta}^{J*}$ at the $k$-th iteration, $\alpha_{k}^{J}$ denotes the $k$-th iteration step size with respect to $\boldsymbol{\theta}_{k}^{J}$, and $\nabla_{\boldsymbol{\theta}^{J}}{\mathbf{L}}_{J}(\boldsymbol{\theta}_{k}^{J})=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}\nabla_{\boldsymbol{\theta}^{J}}\delta_{i}^{TD}(\boldsymbol{\theta}_{k}^{J})$.
Policy update (actor). Since this paper focuses on finding the closed-loop optimal controller by simultaneously tuning the parameters $\boldsymbol{\theta}^{f}$, $\boldsymbol{\theta}^{J}$, and $\boldsymbol{\theta}^{\mu}$. After updating $\boldsymbol{\theta}_{k}^{f}$ and $\boldsymbol{\theta}_{k}^{J}$ following (12) and following (14), respectively, $\boldsymbol{\theta}_{k}^{\mu}$ is updated using the approximated policy gradient described as follows:
$\boldsymbol{\theta}_{k+1}^{\mu}=\boldsymbol{\theta}_{k}^{\mu}-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu}),\quad\boldsymbol{\theta}_{0}^{\mu}\ \text{given},$ (15)
where
$\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}=\frac{1}{N}\sum_{i\in\mathcal{I}_{D}}\Big{(}\nabla_{\boldsymbol{\mu}}c(\boldsymbol{x}_{i},\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu}))\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu})$ (16)
$+\gamma\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{i}^{+},\boldsymbol{\theta}_{k+1}^{J})\nabla_{\boldsymbol{\mu}}\boldsymbol{\hat{x}}_{i}^{+}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu})\Big{)}$
denotes the approximation of policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}$ in (6), obtained by substituting the true system dynamics and $J_{\mu,t^{*}+1}$ in (6) with the DKO dynamics and $V(\boldsymbol{x}(t^{*}+1),\boldsymbol{\theta}_{k+1}^{J})$, respectively, computed using $\mathcal{D}$. Here, for given $\boldsymbol{x}_{i}$, $\boldsymbol{\hat{x}}_{i}^{+}=C_{k}(A_{k}\boldsymbol{g}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k+1}^{f})+B_{k}\boldsymbol{\mu}(\boldsymbol{x}_{i},\boldsymbol{\theta}_{k}^{\mu}))$ represents the one-time-step prediction obtained from DKO and $\nabla_{\boldsymbol{\mu}}\boldsymbol{\hat{x}}_{i}^{+}=C_{k}B_{k}$.
Remark 2
Note that the approximated policy gradient (16) is obtained by differentiating the function $\hat{J}_{\mu}=\frac{1}{N}\sum_{t^{*}\in\mathcal{I}_{D}}\hat{J}_{\mu,t^{*}}$, where $\hat{J}_{\mu,t^{*}}$ denotes an approximation of $J_{\mu,t^{*}}$ in (2) given by
$\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f},\boldsymbol{\theta}^{J},\boldsymbol{\theta}^{\mu})=c(\boldsymbol{x}_{t^{*}},\boldsymbol{\mu}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{\mu}))+\gamma V(\boldsymbol{\hat{x}}(t^{*}+1),\boldsymbol{\theta}^{J})$ (17)
and is constructed in analogy to the expanded form of $J_{\mu,t^{*}}$, given by $J_{\mu,t^{*}}=c(\boldsymbol{x}_{t^{*}},\boldsymbol{\mu}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{\mu}))+\gamma J_{\mu,t^{*}+1}$. If the system dynamics are accurately approximated through DKO, that is, for any given $\boldsymbol{x}_{t^{*}}$ and $\boldsymbol{u}_{t^{*}}$, it holds that $\boldsymbol{x}(t^{*}+1)=C^{*}(A^{*}\boldsymbol{g}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{f*})+B^{*}\boldsymbol{u}_{t^{*}})$. Additionally, if the cost function approximation is effectively learned, i.e., $J_{\mu,t^{*}}=V(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{J*})$. Consequently, the policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}J_{\mu,t^{*}}(\boldsymbol{\theta}^{\mu})$ can be represented by $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu})$. It is worth noting that the computation of $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}$ relies on one-time-step predictions from the DKO dynamics, thereby avoiding the potential accumulation of prediction errors and reducing the computational burden associated with state propagation.
The structure of the proposed policy gradient approximation framework is depicted in Fig. 1,
Figure 1: PGDK framework for policy gradient estimation.
and an offline algorithm to implement this framework is detailed in Algorithm 1.
1 Dataset $\mathcal{D}$. Initialize $\boldsymbol{\theta}_{0}^{f}\in\mathbb{R}^{p}$, $\boldsymbol{\theta}_{0}^{J}\in\mathbb{R}^{s}$, and $\boldsymbol{\theta}_{0}^{\mu}\in\mathbb{R}^{q}$ for $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f})$, $V(\cdot,\boldsymbol{\theta}^{J})$, and $\boldsymbol{\mu}(\cdot,\boldsymbol{\theta}^{\mu})$, respectively. Set step size sequences $\{\alpha_{k}^{f}\}_{k=0}^{K}$, $\{\alpha_{k}^{J}\}_{k=0}^{K}$, $\{\alpha_{k}^{\mu}\}_{k=0}^{K}$, and discount factor $\gamma$.
2 for $k=0,1,\cdots,K$ do
3       Compute $A_{k}$, $B_{k}$ and $C_{k}$ following (13).
4       Update $\boldsymbol{\theta}_{k}^{f}$ and $\boldsymbol{\theta}_{k}^{J}$ following (12) and (14), respectively.
5       Update $\boldsymbol{\theta}_{k}^{\mu}$ following (15).
6 end for
Algorithm 1 Offline policy gradient based on deep Koopman representation (Offline-PGDK)
Note that in practical scenarios, the offline approach presents challenges, particularly when $N$ is large, making dataset acquisition and gradient computation computationally intensive. Moreover, discrepancies between the current policy and the behavior policy employed during offline data collection may impede the identification of the closed-loop optimal controller, as noted in [38]. To address these limitations arising from training data constraints and enhance implementation efficiency, we present an online PGDK algorithm in Section 4.
3.3 Analysis
In this subsection, we conduct a convergence analysis of the proposed framework with respect to the policy parameter $\boldsymbol{\theta}^{\mu}$ based on the given dataset $\mathcal{D}$. To proceed, we formally define the optimal solutions that this study seeks to determine, as presented in the following definition:
Definition 1 (Optimal solutions)
Parameters $\boldsymbol{\theta}^{f*}\in\mathbb{R}^{p}$, $\boldsymbol{\theta}^{J*}\in\mathbb{R}^{s}$, and $\boldsymbol{\theta}^{\mu*}\in\mathbb{R}^{q}$ are optimal solutions if ${\mathbf{L}}_{f}(A^{*},B^{*},C^{*},\boldsymbol{\theta}^{f*})\leq{\mathbf{L}}_{f}(A,B,C,\boldsymbol{\theta}^{f})$, ${\mathbf{L}}_{J}(\boldsymbol{\theta}^{J*})\leq{\mathbf{L}}_{J}(\boldsymbol{\theta}^{J})$, $\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})\leq\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu})$, $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A^{*},B^{*},C^{*},\boldsymbol{\theta}^{f*})=0$, $\nabla_{\boldsymbol{\theta}^{J}}{\mathbf{L}}_{J}(\boldsymbol{\theta}^{J*})=0$, $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})=0$, where $A^{*},B^{*}$ and $C^{*}$ are constant matrices assumed to have full row rank, computed according to (13) with $\boldsymbol{\theta}_{k}^{f}$ replaced by $\boldsymbol{\theta}^{f*}$.
For analytical tractability, we impose the following technical assumption:
Assumption 1
The functions $\boldsymbol{g}(\boldsymbol{x},\boldsymbol{\theta}^{f})$ and $V(\boldsymbol{x},\boldsymbol{\theta}^{J})$ are modeled using linear function approximators, that is, $\boldsymbol{g}(\boldsymbol{x},\boldsymbol{\theta}^{f})=\boldsymbol{\phi}_{f}(\boldsymbol{x})\boldsymbol{\theta}^{f}$ and $V(\boldsymbol{x},\boldsymbol{\theta}^{J})=\boldsymbol{\phi}_{J}(\boldsymbol{x})^{\prime}\boldsymbol{\theta}^{J}$, where $\boldsymbol{\phi}_{f}(\boldsymbol{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r\times p}$ and $\boldsymbol{\phi}_{J}(\boldsymbol{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{s}$ are manually chosen basis functions satisfying $\parallel\boldsymbol{\phi}_{f}(\boldsymbol{x})\parallel\leq 1$ and $\parallel\boldsymbol{\phi}_{J}(\boldsymbol{x})\parallel\leq 1$.
Given that the proposed framework seeks to minimize (4) by simultaneously tuning $\boldsymbol{\theta}^{f}$, $\boldsymbol{\theta}^{J}$, and $\boldsymbol{\theta}^{\mu}$, where $\boldsymbol{\theta}^{f}$ and $\boldsymbol{\theta}^{J}$ are updated independently while $\boldsymbol{\theta}^{\mu}$ is updated based on $\boldsymbol{\theta}^{f}$ and $\boldsymbol{\theta}^{J}$, it is essential to first analyze the convergence properties of the parameters $\boldsymbol{\theta}_{k+1}^{f}$ and $\boldsymbol{\theta}_{k+1}^{J}$. Prior to presenting the main results regarding the approximated policy gradient, we establish the following key result related to the dynamics approximation:
Lemma 1
If Assumption 1 holds and $\boldsymbol{\theta}^{f}$ is updated following (12), for constant step size satisfying $0<\alpha_{f}<\frac{1}{2L_{f1}}$, one has $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}=\frac{\alpha_{f}L_{f2}}{2L_{f1}}$, where $L_{f1}$ is a constant, and $L_{f2}$ denotes a constant determined by the residual vectors associated with the least squares solutions in (13), as detailed in the Appendix. To obtain $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}=0$, one needs to appropriately choose a diminishing step size satisfying $0<\alpha_{k}^{f}<\frac{1}{2L_{f1}}$, with $\sum_{k=0}^{\infty}\alpha_{k}^{f}=\infty$ and $\sum_{k=0}^{\infty}(\alpha_{k}^{f})^{2}<\infty$.
The proof of Lemma 1 is provided in the Appendix. Based on Lemma 1, we establish the following corollary:
Corollary 1
Under the same assumptions as in Lemma 1, choosing the step size $\alpha_{k}^{f}=\frac{\beta_{f}}{2+k}$ with $\beta_{f}=\frac{1}{L_{f1}}$ yields
$\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\leq\frac{\nu_{f}}{2+k},$ (18)
where $\nu_{f}=\max\{\beta_{f}^{2}L_{f2},\quad 2\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\}$.
Remark 3
Since $L_{f2}$ in (18) is derived from the residual vectors from the least squares solutions in (13), it generally results in a smaller value of $L_{f2}$ compared to using the arbitrarily assigning the constant matrices $A$, $B$, and $C$. This reduction in $L_{f2}$ consequently leads to a tighter upper bound in (18), thereby expediting convergence of $\boldsymbol{\theta}_{k}^{f}\rightarrow\boldsymbol{\theta}^{f*}$.
As TD error learning is a well-established method, we refer to the seminal work [39] for a detailed theoretical analysis. In particular, Theorem $1$ in [39] proves that under certain assumptions, when $\boldsymbol{\theta}_{k}^{J}$ is updated according to (14), it holds that $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{J}-\boldsymbol{\theta}^{J*}\parallel^{2}=0$. A detailed explanation of this result is provided in the Appendix.
To examine the convergence behavior of $\boldsymbol{\theta}_{k}^{\mu}$, we first establish the following consequence to demonstrate the Lipschitz continuity of the introduced $\hat{J}_{\mu,t^{*}}$ in (17):
Lemma 2
If Assumption 1 holds, then the introduced $\hat{J}_{\mu,t^{*}}$ in (17) and its gradient $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}$ are also Lipschitz continuous with respect to $\boldsymbol{\theta}^{\mu}$, with Lipschitz constants $L_{J\theta}$ and $L_{J\theta\theta}$, respectively, as detailed in the Appendix.
The proof of Lemma 2 is provided in the Appendix. Building upon the preceding results, as $\boldsymbol{\theta}_{k}^{f}\rightarrow\boldsymbol{\theta}^{f*}$ and $\boldsymbol{\theta}_{k}^{J}\rightarrow\boldsymbol{\theta}^{J*}$, we have the following convergence properties of the approximated policy gradient:
Theorem 1
Under the same assumptions as in Lemma 1 and Lemma 2, if the parameters $\boldsymbol{\theta}_{k}^{f}$, $\boldsymbol{\theta}_{k}^{J}$, and $\boldsymbol{\theta}_{k}^{\mu}$ are updated according to (12), (14), and (15), respectively, using the step sizes specified in Corollary 1 and Lemma 3. To attain $\lim_{k\rightarrow\infty}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}=0$, it is necessary to use diminishing step size $\alpha_{k}^{\mu}$ satisfying $\sum_{k=0}^{\infty}\alpha_{k}^{\mu}=\infty$ and $\sum_{k=0}^{\infty}(\alpha_{k}^{\mu})^{2}<\infty$.
The proof of Theorem 1 is provided in the Appendix. As a direct consequence of Theorem 1, we state the following corollary:
Corollary 2
If the assumptions in Theorem 1 hold, and the step sizes are chosen as $\alpha_{k}^{f}=\frac{\beta_{f}}{2+k}$ from Corollary 1, $\alpha_{J}=(1-\gamma)/4$ from Lemma 3, and $\alpha_{k}^{\mu}=(k+1)^{-1/4}$, then the estimated policy gradient in (16) has the following property:
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$
$\leq$ $L_{a}K^{-1/4}+L_{b}K^{-3/4}+L_{c}K^{-5/4},$
where $L_{b}=\frac{3}{4}(\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})+C_{1}(\frac{2}{1-\sqrt{1-\nu_{J}}}+(1-\nu_{J})^{K/2})\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel)$, and constants $L_{a}$, $L_{c}$ and $C_{1}$ are specified in the Appendix, while $10<\nu_{J}<$ 1 is a constant defined in Lemma 3.
Remark 4
Theorem 1 and Corollary 2 establish the convergence properties of the estimated policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}$ in (16) under appropriately chosen step sizes. It states that, under certain assumptions, if the dynamics and cost function are successfully learned, that is, $\boldsymbol{x}(t+1)=C^{*}(A^{*}\boldsymbol{g}(\boldsymbol{x}(t),\boldsymbol{\theta}^{f*})+B^{*}\boldsymbol{u}(t))$ and $J_{\mu,t^{*}}=V(\boldsymbol{x}(t^{*}),\boldsymbol{\theta}^{J*})$, then the proposed framework can achieve a globally optimal closed-loop controller, provided that the functions $c$ and $V$ are convex. In the case where these functions are non-convex, the framework is capable of obtaining a locally optimal closed-loop controller.
In the following section, we validate the theoretical results presented above and demonstrate the performance of the proposed framework through numerical simulations.
4 Numerical Simulations
In this section, we first present an online algorithm to implement the proposed framework efficiently. Then, we evaluate the performance of both the offline and online algorithms through numerical simulations.
4.1 Online Implementation
We now present an online algorithm for the efficient implementation of the proposed policy gradient estimation framework, which employs the widely used mini-batch gradient descent method [40] to update the parameters $\boldsymbol{\theta}_{k}^{f}$, $\boldsymbol{\theta}_{k}^{J}$, and $\boldsymbol{\theta}_{k}^{\mu}$, and is detailed below.
Data generation. To ensure the sufficient exploration of the system state space $\mathcal{X}$, we adopt the widely-used method from [8], where an exploration policy is implemented by introducing noise $W(t)\in\mathcal{W}\subset\mathbb{R}^{m}$, sampled from a noise process, into the policy in (3) at each time step $t$ while updating $\boldsymbol{\theta}_{k}^{\mu}$ following (5). Common choices for $W(t)$ include the Ornstein-Uhlenbeck process [41] or modeling it as a Gaussian distribution. Specifically, let $t_{i}$ denote the system time at the $i$-th observation. Given system state $\boldsymbol{x}(t_{i})=\boldsymbol{x}_{t_{i}}$, by applying
$\boldsymbol{\bar{u}}(t_{i})=\boldsymbol{\mu}(\boldsymbol{x}(t_{i}),\boldsymbol{\theta}_{k}^{\mu})+\sigma(t_{i})W(t_{i})$ (19)
into the unknown dynamics in (1), the system state at the next time step, $\boldsymbol{x}(t_{i}+1)$ is observed, where $\sigma(t_{i})\geq 0$ is a time-decay function designed to ensure that the effect of $W(t_{i})$ on $\boldsymbol{\mu}$ gradually decreases as $\boldsymbol{\theta}_{k}^{\mu}$ approaches the optimal value $\boldsymbol{\theta}^{\mu*}$. To differentiate between system state-input variables and observed constant system state-input data pairs, we denote $(\boldsymbol{x}_{t_{i}},\boldsymbol{\bar{u}}_{t_{i}},\boldsymbol{x}_{t_{i}}^{+})$ as the observed constant tuple corresponding to $(\boldsymbol{x}(t_{i}),\boldsymbol{\bar{u}}(t_{i}),\boldsymbol{x}(t_{i}+1))$. For notational brevity, we write $(\boldsymbol{x}_{i},\boldsymbol{u}_{i},\boldsymbol{x}_{i}^{+})\coloneqq(\boldsymbol{x}_{t_{i}},\boldsymbol{\bar{u}}_{t_{i}},\boldsymbol{x}_{t_{i}}^{+})$ throughout the remainder of this paper. The observed tuples $(\boldsymbol{x}_{i},\boldsymbol{u}_{i},\boldsymbol{x}_{i}^{+})$ are stored in a finite-sized data memory $\mathcal{D}_{i}$, which has a fixed maximum capacity $N$, and is defined as follows:
$\mathcal{D}_{i}=\{(\boldsymbol{x}_{1},\boldsymbol{u}_{1},\boldsymbol{x}_{1}^{+}),(\boldsymbol{x}_{2},\boldsymbol{u}_{2},\boldsymbol{x}_{2}^{+}),\cdots,(\boldsymbol{x}_{i},\boldsymbol{u}_{i},\boldsymbol{x}_{i}^{+})\},i\leq N.$
For $i>N$, the earliest $i-N$ observed tuples are discarded to accommodate new data. The value of $N$ is typically chosen to be large. At any iteration $k$, a data batch of $\tilde{N}$ tuples ($\tilde{N}\ll N$) is sampled from $\mathcal{D}_{i}$, with its index set $\mathcal{I}_{k}$ drawn uniformly from all subsets of $\{1,2,\cdots,i\}$ with size $\tilde{N}$, provided that $i\geq\tilde{N}$. This batch is then used to estimate the gradients in (12), (14), and (16). Note that, under the online setting, data collection and parameter updates are performed concurrently.
Parameters tuning using sampled data batches. Given the above sampled data batch from $\mathcal{D}_{i}$ at any iteration $k$, we construct the following data matrices, which are exclusively utilized within the scope of this subsection:
${\mathbf{X}}_{k}$ $=[\boldsymbol{x}_{1},\boldsymbol{x}_{2},\cdots,\boldsymbol{x}_{\tilde{N}}]\in\mathbb{R}^{n\times\tilde{N}},$ (20)
$\mathbf{U}_{k}$ $=[\boldsymbol{u}_{1},\boldsymbol{u}_{2},\cdots,\boldsymbol{u}_{\tilde{N}}]\in\mathbb{R}^{m\times\tilde{N}},$
$\bar{{\mathbf{X}}}_{k}$ $=[\boldsymbol{x}_{1}^{+},\boldsymbol{x}_{2}^{+},\cdots,\boldsymbol{x}_{\tilde{N}}^{+}]\in\mathbb{R}^{n\times\tilde{N}},$
${\mathbf{G}}_{k}$ $=[\boldsymbol{g}(\boldsymbol{x}_{1},\boldsymbol{\theta}_{k}^{f}),\boldsymbol{g}(\boldsymbol{x}_{2},\boldsymbol{\theta}_{k}^{f}),\cdots,\boldsymbol{g}(\boldsymbol{x}_{\tilde{N}},\boldsymbol{\theta}_{k}^{f})]\in\mathbb{R}^{r\times\tilde{N}},$
$\bar{{\mathbf{G}}}_{k}$ $=[\boldsymbol{g}(\boldsymbol{x}_{1}^{+},\boldsymbol{\theta}_{k}^{f}),\boldsymbol{g}(\boldsymbol{x}_{2}^{+},\boldsymbol{\theta}_{k}^{f}),\cdots,\boldsymbol{g}(\boldsymbol{x}_{\tilde{N}}^{+},\boldsymbol{\theta}_{k}^{f})]\in\mathbb{R}^{r\times\tilde{N}},$
where $(\boldsymbol{x}_{\tilde{N}},\boldsymbol{u}_{\tilde{N}},\boldsymbol{x}_{\tilde{N}}^{+})$ represents the $\tilde{N}$-th data tuple in the sampled batch, as used exclusively in (20) for notational convenience. This notation is employed because the objective is to find the mapping between $\boldsymbol{x}_{t},\boldsymbol{u}_{t}$ and $\boldsymbol{x}_{t}^{+}$ without regard to the specific time at which it is recorded.
If the matrices $\bar{{\mathbf{G}}}_{k}\in\mathbb{R}^{r\times\tilde{N}}$ and $\begin{bmatrix}{\mathbf{G}}_{k}\\ \mathbf{U}_{k}\end{bmatrix}\in\mathbb{R}^{(r+m)\times\tilde{N}}$ are with full row rank, then the parameter $\boldsymbol{\theta}_{k}^{f}$ is updated in the same manner as (12) with its gradient computed using the sampled data batch, where $[A_{k},B_{k}]$ and $C_{k}$ are computed according to (13) using the data matrices defined in (20).
After updating $\boldsymbol{\theta}_{k}^{f}$, the parameters $\boldsymbol{\theta}_{k}^{J}$ and $\boldsymbol{\theta}_{k}^{\mu}$ are updated in a manner analogous to (14) and (15), respectively, using gradients computed over the sampled data batch. In summary, we introduce an online PGDK in Algorithm 2.
1 Initialize $\boldsymbol{\theta}_{0}^{f}\in\mathbb{R}^{p}$, $\boldsymbol{\theta}_{0}^{J}\in\mathbb{R}^{s}$, and $\boldsymbol{\theta}_{0}^{\mu}\in\mathbb{R}^{q}$ for $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f})$, $V(\cdot,\boldsymbol{\theta}^{J})$, and $\boldsymbol{\mu}(\cdot,\boldsymbol{\theta}^{\mu})$, respectively. Set the iteration index $k=0$ and step size sequences $\{\alpha_{k}^{f}\}_{k=0}^{K}$, $\{\alpha_{k}^{J}\}_{k=0}^{K}$, $\{\alpha_{k}^{\mu}\}_{k=0}^{K}$. Specify discount factor $\gamma$, number of episodes $E$, task horizon $T$, batch size $\tilde{N}$, time-decay function $\sigma(t)$, and data memory.
2 for $\textnormal{episode}=1,2,\cdots,E$ do
3      Reset the initial state $\boldsymbol{x}_{0}$ and initialize the noise process $W$.
4       for $i=0,1,\cdots,T$ do
5             Execute the control input $\boldsymbol{\bar{u}}(t_{i})=\boldsymbol{\mu}(\boldsymbol{x}(t_{i}),\boldsymbol{\theta}_{k}^{\mu})+\sigma(t_{i})W(t_{i})$, observe the resulting $\boldsymbol{x}(t_{i}+1)$, and store the observed tuple $(\boldsymbol{x}_{t_{i}},\boldsymbol{\bar{u}}_{t_{i}},\boldsymbol{x}_{t_{i}+1})$ in the data memory.
6             Sample $\tilde{N}$ data tuples uniformly from the data memory.
7             Update $\boldsymbol{\theta}_{k}^{f}$ analogous to (12) using the batch gradient, where $A_{k}$, $B_{k}$, and $C_{k}$ are computed following (13) using (20).
8             Update $\boldsymbol{\theta}_{k}^{J}$ and $\boldsymbol{\theta}_{k}^{\mu}$ in a manner analogous to (14) and (15), respectively, using the gradient computed over the sampled data batch.
9             Set $k=k+1$.
10       end for
11
12 end for
Algorithm 2 Online policy gradient based on deep Koopman representation (Online-PGDK)
4.2 Numerical Simulations
In this subsection, we apply the proposed algorithms to derive closed-loop optimal controllers for both a linear time-invariant (LTI) system and a nonlinear inverted pendulum example. Following this, we compare the performance of the proposed methods with related baseline algorithms, highlighting key differences in control strategies and convergence behavior. Finally, we provide a detailed performance analysis, offering insights into the advantages and limitations of the proposed PGDK framework in various settings.
4.2.1 LTI System
Consider the following LTI dynamics:
$\boldsymbol{x}(t+1)=\begin{bmatrix}&0.5&0.5\\ &0&1\end{bmatrix}\boldsymbol{x}(t)+\begin{bmatrix}0\\ 1\end{bmatrix}u(t),\quad\boldsymbol{x}(0)=\boldsymbol{x}_{0},$
where $[-5,-5]^{\prime}\leq\boldsymbol{x}(t)\leq[5,5]^{\prime}$ and $-1\leq u(t)\leq 1$ denote the system state and control input at time $t$, respectively.
Setup. For each simulation episode, the initial state $\boldsymbol{x}_{0}$ is uniformly sampled from the interval $[-0.1,-0.1]^{\prime}$ and $[0.1,0.1]^{\prime}$, and each episode is terminated either when $t\geq 50$. The objective of this example is to design a closed-loop optimal controller for driving the system state toward the goal state $\boldsymbol{x}_{goal}=[1,1]^{\prime}$ starting from any given initial state, for which we design the following stage cost function:
$c(\boldsymbol{x}(t),u(t))=(\boldsymbol{x}(t)-\boldsymbol{x}_{goal})^{\prime}(\boldsymbol{x}(t)-\boldsymbol{x}_{goal})+0.001u(t)^{2}.$
To conduct the simulation, we construct the deep Koopman basis function $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f}):\mathbb{R}^{2}\rightarrow\mathbb{R}^{4}$ to lift the original system states into a higher-dimensional space, which is applied in both the offline and online PGDK algorithms. We use the $ReLU$ activation function to build the $2$-layer architecture DNNs $\boldsymbol{g}$, $V$, and $\boldsymbol{\mu}$. In particular, we execute the online PGDK algorithm first to generate a data memory, which is then used as the training dataset for the offline PGDK algorithm. All DNNs are trained using the Adam optimizer [42].
Evaluation. In this simulation, we are interested in evaluating how the proposed offline and online PGDK perform on a simple control task compared to the classical optimal control technique, the linear quadratic regulator (LQR), which has access to the exact system dynamics and employs the same stage cost function. To assess the performance of the algorithms, we compare them by presenting both the simulation costs and the optimal trajectories achieved by each method, using identical initial states to ensure a fair comparison. To mitigate the influence of randomness in the DNN training, the entire experiment is repeated over $5$ independent trials.
Analysis. As shown in the first subplot of Fig. 2, the online PGDK algorithm demonstrates a convergence rate close to that of the offline PGDK approach with respect to the number of training episodes. Furthermore, the second subplot of Fig. 2 shows that the closed-loop controller learned via online PGDK attains stage costs that are consistently closer to those of the benchmark LQR controller relative to the offline PGDK implementation, when evaluated from identical initial states. This improvement in stage cost can be attributed to the fact that the offline PGDK is more prone to converging to local minima. In contrast, the online PGDK, which updates gradients using mini-batches sampled from a continually refreshed replay buffer, is more likely to escape local minima. Notably, the error bars in the second subplot indicate that the best-performing instances of the proposed PGDK algorithms across $5$ trials attain performance close to the LQR controller, underscoring the potential of the PGDK framework to achieve control performance close to classical optimal control methods when properly trained.
Figure 2: Learning and testing stage cost using the same initial states. Here, $\bar{c}$ denotes the averaged stage cost over each episode to account for the variance of different initial states. The solid line represents the mean stage cost across $5$ experiment trials, while the shaded region and error bars indicate the standard deviation.
Fig. 3 illustrates an example trajectory comparison between the proposed PGDK algorithms and the LQR controller. It can be observed that both online and offline PGDK algorithms tend to apply more aggressive control inputs than the LQR controller in this LTI system scenario. This behavior results in faster convergence toward the goal state, albeit at the cost of potentially higher control energy. The LQR controller, on the other hand, follows a smoother trajectory with more conservative control inputs, reflecting its optimality under a quadratic cost formulation.
Figure 3: Trajectories from PGDK and LQR.
4.2.2 Simulated Inverted Pendulum
We now illustrate the performance of the proposed algorithms using the nonlinear inverted pendulum example, of which the system dynamics are given by
$\begin{cases}\theta(t+1)=\theta(t)+\dot{\theta}(t+1)\Delta t,\\ \dot{\theta}(t+1)=\dot{\theta}(t)+(\frac{-3g\sin(\theta(t)+\pi)}{2l}+\frac{3u(t)}{ml^{2}})\Delta t,\end{cases}$ (21)
where $g$, $m$, and $l$ represent the acceleration due to gravity, the mass of the pendulum, and the length of the pendulum, respectively. Here, the system state is defined as $\boldsymbol{x}(t)=[\theta(t),\dot{\theta}(t)]^{\prime}$ with the lower bound and upper bound of $[-\pi,-8]^{\prime}$ and $[\pi,8]^{\prime}$, respectively, and $u(t)$ denotes the continuous scalar torque input, constrained between $-2$ and $2$.
Setup. For each simulation episode, the initial state $\boldsymbol{x}(0)$ is sampled uniformly from the interval $[-\pi,-1]^{\prime}$ and $[\pi,1]^{\prime}$ and an episode terminates if $t>200$. The goal is to balance the pendulum to the upright position, represented by $\boldsymbol{x}_{goal}=[0,0]^{\prime}$. To achieve this, the following stage cost function is defined:
$c(\boldsymbol{x}(t),\boldsymbol{u}(t))=\theta(t)^{2}+0.1\dot{\theta}(t)^{2}+0.001u(t)^{2}.$
To conduct the simulation, we set $\Delta t=0.02s$, and lift the system states using the deep Koopman basis function $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f}):\mathbb{R}^{2}\rightarrow\mathbb{R}^{8}$, which is utilized in both the offline and online PGDK algorithms. We use the DNNs with the same structures and activation functions as in the LTI example, except an added layer in the policy $\boldsymbol{\mu}$, which uses a $Tanh()$ activation function to enhance expressiveness for the nonlinear dynamics. The online PGDK algorithm is executed first, and the data memory generated during this process is retained and subsequently used as the training dataset for the offline PGDK implementation. We use the Adam optimizer for DNN training.
Evaluation. This simulation aims to evaluate the convergence rate of the proposed online PGDK algorithm in comparison with existing on-policy reinforcement learning (RL) methods, as well as to assess the optimality of the resulting closed-loop controller relative to model-based optimal control strategies. Following this, to assess the performance of the proposed online PGDK, we initially compare it with a representative model-free RL algorithm, the deep deterministic policy gradient (DDPG) [8], to highlight data efficiency. Subsequently, we evaluate the online PGDK alongside two additional approaches: MPC with a known exact dynamics model and MPC with a learned deep Koopman model (DKMPC) [31]. To demonstrate the optimality of the policy learned by PGDK, we compare the trial cost, $J=\sum_{t=0}^{200}c(\boldsymbol{x}(t),\boldsymbol{u}(t))$, across PGDK, MPC, and DKMPC. The comparison uses identical stage cost functions and the following seven initial states: $(\theta_{0}=\pi/12,\dot{\theta}_{0}=-1)$, $(\theta_{0}=-\pi/12,\dot{\theta}_{0}=-1)$, $(\theta_{0}=\pi/4,\dot{\theta}_{0}=1)$, $(\theta_{0}=-\pi/4,\dot{\theta}_{0}=1)$, $(\theta_{0}=\pi/2,\dot{\theta}_{0}=0)$, $(\theta_{0}=-\pi/2,\dot{\theta}_{0}=0)$, and $(\theta_{0}=\pi,\dot{\theta}_{0}=0)$. The prediction horizons for MPC and DKMPC are set to $80$ and $6$, respectively.
Analysis. As demonstrated in Fig. 4, both offline and online PGDK methods exhibit asymptotic convergence, consistent with theoretical expectations. Notably, offline PGDK reaches policy convergence in fewer iterations since it computes gradients over the entire dataset, while the online PGDK estimates gradients using sampled data batches. However, in practical applications, offline PGDK typically demands more computational time due to the necessity of processing the full dataset for parameter updates.
Figure 4: Learning losses of the online (blue) and offline (black) PGDK, where $L_{f}$, $L_{J}$, and $\hat{J}_{\mu}$ represent the loss functions computed over the entire dataset at iteration $k$, while $L_{k}^{f}$, $L_{k}^{J}$, and $\hat{J}_{k}^{\mu}$ denote their corresponding values computed using the sampled data batches.
As can be seen in Fig. 5, for the inverted pendulum example, the proposed online PGDK can reach DDPG’s performance but with better data efficiency since it requires fewer episodes ($10$ episodes) to find the optimal policy compared to the DDPG ($60$ episodes).
Figure 5: Learning stage cost of online PGDK and DDPG, where the dashed line denotes the average stage cost of each episode to account for the variability in initial states, and the solid lines denote the cumulative average stage cost.
Finally, Fig. 6 shows the performance of the proposed online PGDK compared to the model-based MPC approach with the same initial states. As shown by the testing trial costs, online PGDK achieves performance close to MPC that has access to the exact system dynamics. Furthermore, online PGDK outperforms MPC when the latter relies on trajectory propagation via a learned deep Koopman model, which is susceptible to cumulative model approximation errors.
Figure 6: Testing cost of online PGDK, MPC, and DKMPC.
5 Concluding Remarks
In this paper, we have described a data-driven policy gradient approximation framework, termed policy gradient with deep Koopman representation (PGDK), which is designed to derive a closed-loop optimal controller for systems with unknown dynamics. Additionally, we have introduced both an offline and an online algorithm to implement the proposed framework. The key contribution of PGDK lies in its integration of deep Koopman learning within the concepts of actor-critic methods to approximate the policy gradient in (6) with (16), enabling the simultaneous approximation of system dynamics, cost functions, and optimal policies. This integration enhances data efficiency and accelerates convergence compared to conventional deterministic policy gradient methods. Furthermore, we present a theoretical convergence analysis of the proposed framework, supported by numerical simulations. Finally, through experiments on an LTI system and a pendulum example, we demonstrate that the closed-loop optimal controller derived via the proposed algorithms for an unknown dynamical system achieves performance close to that of an optimal controller designed with known system dynamics within finite tuning iterations.
6 Appendix
In this section, we present the proofs of the theoretical results discussed in Section 3.3. Before delving into the main proofs, we introduce the following supporting lemma, which pertains to the convergence analysis of the TD error:
Lemma 3
[Theorem $1$, [39]] Suppose the DNN $V(\boldsymbol{x},\boldsymbol{\theta}^{J})$ is modeled using a linear function approximation as $V(\boldsymbol{x},\boldsymbol{\theta}^{J})=\boldsymbol{\phi}_{J}(\boldsymbol{x})^{\prime}\boldsymbol{\theta}^{J}$, where $\parallel\boldsymbol{\phi}_{J}(\boldsymbol{x})\parallel\leq 1$. Additionally, it is assumed that the system dynamics in (1) following the policy $\boldsymbol{\mu}$ is ergodic with a unique stationary distribution $\pi(\boldsymbol{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$. Here, for any two states $\boldsymbol{x}_{1},\boldsymbol{x}_{2}:\pi(\boldsymbol{x}_{2})=\lim_{t\rightarrow\infty}\mathbb{P}(\boldsymbol{x}_{t}=\boldsymbol{x}_{2}|\boldsymbol{x}_{0}=\boldsymbol{x}_{1})$. For $\Sigma=\sum_{\boldsymbol{x}\in\mathcal{X}}\pi(\boldsymbol{x})\boldsymbol{\phi}_{J}(\boldsymbol{x})\boldsymbol{\phi}_{J}(\boldsymbol{x})^{\prime}$, let $\omega_{J}$ be the smallest eigenvalue of $\Sigma$. Given $\omega_{J}>0$ and any constant step size $\alpha_{J}=(1-\gamma)/4$, let $\nu_{J}=(1-\gamma)^{2}\omega_{J}/4$ by following the updating rule described in (14) one has
$\parallel\boldsymbol{\theta}_{k}^{J}-\boldsymbol{\theta}^{J*}\parallel^{2}\leq(1-\nu_{J})^{k}\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel^{2}.$
6.1 Proof of Lemma 1
In this subsection, we present the convergence proof for the proposed updating rule in (12). To proceed, we first recall the loss function ${\mathbf{L}}_{f}(A,B,C,\boldsymbol{\theta}^{f})$ from (10) and data matrices from (11). Given that the DNN $\boldsymbol{g}(\boldsymbol{x},\boldsymbol{\theta}^{f})$ is constructed using a linear function approximator of the form: $\boldsymbol{g}(\boldsymbol{x},\boldsymbol{\theta}^{f})=\boldsymbol{\phi}_{f}(\boldsymbol{x})\boldsymbol{\theta}^{f}$, where $\parallel\boldsymbol{\phi}_{f}(\boldsymbol{x})\parallel\leq 1$, we reformulate ${\mathbf{L}}_{f}(A,B,C,\boldsymbol{\theta}^{f})$ into a compact form using the definition of the Frobenius norm, given by
${\mathbf{L}}_{f}(A,B,C,\boldsymbol{\theta}^{f})$
$=$ $\frac{1}{2N}(\parallel\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}^{f}\!\!-\!\!A\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}^{f}\!\!-\!\!B\mathbf{U}\parallel_{F}^{2}+\parallel\mathbf{\bar{X}}\!-\!C\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}^{f}\parallel_{F}^{2}),$
where $\boldsymbol{\bar{\Phi}}_{f}=[\boldsymbol{\phi}_{f}(\boldsymbol{x}_{1}^{+}),\boldsymbol{\phi}_{f}(\boldsymbol{x}_{2}^{+}),\cdots,\boldsymbol{\phi}_{f}(\boldsymbol{x}_{N}^{+})]\in\mathbb{R}^{r\times Np}$ and $\boldsymbol{\Phi}_{f}=[\boldsymbol{\phi}_{f}(\boldsymbol{x}_{1}),\boldsymbol{\phi}_{f}(\boldsymbol{x}_{2}),\cdots,\boldsymbol{\phi}_{f}(\boldsymbol{x}_{N})]\in\mathbb{R}^{r\times Np}$ are constant matrices. Here, we denote $\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}^{f}=[\boldsymbol{\phi}_{f}(\boldsymbol{x}_{1})\boldsymbol{\theta}^{f},\boldsymbol{\phi}_{f}(\boldsymbol{x}_{2})\boldsymbol{\theta}^{f},\cdots,\boldsymbol{\phi}_{f}(\boldsymbol{x}_{N})\boldsymbol{\theta}^{f}]\in\mathbb{R}^{r\times N}$ for the sake of notational simplicity. Note that $\boldsymbol{\Phi}_{f}\cdot(\boldsymbol{\theta}_{1}^{f}+\boldsymbol{\theta}_{2}^{f})=\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}_{1}^{f}+\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}_{2}^{f}$, $\parallel\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}^{f}\parallel\leq\parallel\boldsymbol{\Phi}_{f}\parallel\parallel\boldsymbol{\theta}^{f}\parallel$, and $\nabla_{\boldsymbol{\theta}^{f}}\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}^{f}=\boldsymbol{\Phi}_{f}$.
As stated in Definition 1, $\boldsymbol{\theta}^{f*}$ represents the optimal parameter satisfying ${\mathbf{L}}_{f}(A^{*},B^{*},C^{*},\boldsymbol{\theta}^{f*})\leq{\mathbf{L}}_{f}(A,B,C,\boldsymbol{\theta}^{f})$. For brevity, we define $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}=\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f})$. To establish that $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}=0$, we begin by taking the squared norm after subtracting $\boldsymbol{\theta}^{f*}$ on both sides of (12), which results in:
$\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}$ (22)
$=$ $\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}-\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}$
$=$ $\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\!\!-2\alpha_{k}^{f}\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k},\boldsymbol{\theta}_{k}^{f}\!-\!\boldsymbol{\theta}^{f*}\rangle+\!\!\parallel\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}\!\!\!.$
The convergence proof of (22) consists of two main parts. First, we establish an upper bound for $-\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k},\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$. Then, we derive an upper bound for $\parallel\alpha_{k}^{f}\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}$.
To begin, let $\Delta\boldsymbol{\theta}_{k}^{f}=\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}$, we derive a upper bound of $-\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k},\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$ by expanding $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}$, yielding:
$\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f}),\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$ (23)
$=$ $\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f})-\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}^{f*})$
$+\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}^{f*}),\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$
$=$ $\frac{1}{N}(\parallel\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f}\parallel_{F}^{2}+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2})\parallel\Delta\boldsymbol{\theta}_{k}^{f}\parallel^{2}$
$+\frac{1}{2}\Big{(}{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}^{f*})-{\mathbf{L}}_{f}(A_{k},B_{k},C_{k},\boldsymbol{\theta}_{k}^{f})$
$-\frac{1}{N}(\parallel(\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f})\cdot\Delta\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\cdot\Delta\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2})\Big{)}$
$=$ $\frac{3}{2N}(\parallel\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f}\parallel_{F}^{2}+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2})\parallel\Delta\boldsymbol{\theta}_{k}^{f}\parallel^{2}$
$-\frac{1}{N}(\parallel(\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f})\cdot\Delta\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\cdot\Delta\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}),$
where the last equality of (23) is obtained by applying the Taylor expansion. Furthermore, based on (23), we obtain the following result:
$-\langle\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k},\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\rangle$ (24)
$\leq$ $-\frac{1}{2N}(\parallel\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f}\parallel_{F}^{2}+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2})\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}.$
To derive an upper bound for (24), we utilize the cyclic property of the trace operator and apply the Frobenius norm bound from [43]. Specifically, for matrices $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times r}$, the following inequality holds:
$\lambda_{min}(A)Tr(B)\leq Tr(AB)\leq\lambda_{max}(A)Tr(B).$
Thus, we obtain
$\frac{1}{2N}(\parallel\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2}\!+\!\parallel A_{k}\boldsymbol{\Phi}_{f}\parallel_{F}^{2}\!+\!\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2}\!\!-2\langle\boldsymbol{\bar{\Phi}}_{f},\!A_{k}\boldsymbol{\Phi}_{f}\rangle_{F})$ (25)
$\geq$ $\frac{1}{2N}(\omega_{f_{1}}(Np+\!\parallel A_{k}\parallel_{F}^{2}\!+\!\parallel C_{k}\parallel_{F}^{2})\!-\!2\omega_{f_{2}}Tr(A_{k}))\eqqcolon L_{f1},$
where $\omega_{f_{1}}=\min\{\lambda_{min}(\boldsymbol{\bar{\Phi}}_{f}\boldsymbol{\bar{\Phi}}_{f}^{\prime}),\lambda_{min}(\boldsymbol{\Phi}_{f}\boldsymbol{\Phi}_{f}^{\prime})\}$ and $\omega_{f_{2}}=\lambda_{max}(\boldsymbol{\bar{\Phi}}_{f}\boldsymbol{\Phi}_{f}^{\prime})$. Here, a sufficiently large $N$ may be required to ensure $L_{f1}>0$.
To establish an upper bound for $\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}$, we proceed by utilizing the definition of ${\mathbf{L}}_{f}$, leading to the following result:
$\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}$
$=$ $\frac{1}{N}\parallel(\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f})^{\prime}(\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-A\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-B_{k}\mathbf{U})$
$-(C_{k}\boldsymbol{\bar{\Phi}}_{f})^{\prime}(\mathbf{\bar{X}}-C\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f})\parallel_{F}^{2}$
$\leq$ $\frac{2}{N}(\parallel\boldsymbol{\bar{\Phi}}_{f}-A_{k}\boldsymbol{\Phi}_{f}\parallel_{F}^{2}\parallel\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-A_{k}\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-B_{k}\mathbf{U}\parallel_{F}^{2}$
$+\parallel C_{k}\boldsymbol{\bar{\Phi}}_{f}\parallel_{F}^{2}\parallel\bar{{\mathbf{X}}}-C_{k}\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}),$
where $[A_{k}\ B_{k}]$ and $C_{k}$ are obtained via the least squares solutions in (13). Here, given that the system states and control inputs are bounded and $\boldsymbol{g}$ is designed to be Lipschitz continuous, the norms of the matrices $A_{k}$, $B_{k}$, and $C_{k}$ from (13) remain bounded. Specifically, we define these bounds as $\parallel A_{k}\parallel\leq c_{1}$, $\parallel B_{k}\parallel\leq c_{2}$, and $\parallel C_{k}\parallel\leq c_{3}$ throughout this proof. Furthermore, we define the upper bounds of the residual vectors originating from the least squares solutions as [44] $\parallel\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-A_{k}\boldsymbol{\Phi}_{f}\cdot\boldsymbol{\theta}_{k}^{f}-B_{k}\mathbf{U}\parallel_{F}^{2}\leq W_{1}$ and $\parallel\bar{{\mathbf{X}}}-C_{k}\boldsymbol{\bar{\Phi}}_{f}\cdot\boldsymbol{\theta}_{k}^{f}\parallel_{F}^{2}\leq W_{2}$, which leads to following upper bound:
$\parallel\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f,k}\parallel^{2}\leq(2(1+c_{1}^{2})W_{1}+2c_{3}^{2}W_{2})/N\coloneqq L_{f2}.$ (26)
Finally, applying the derived bounds $L_{f1}$ from (25) and $L_{f2}$ from (26) to (22), we obtain:
$\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\leq(1-2\alpha_{k}^{f}L_{f1})\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}+(\alpha_{k}^{f})^{2}L_{f2}.$ (27)
Here, to achieve $10<1-2\alpha_{k}^{f}L_{f1}<$ 1, one needs to choose step size $0<\alpha_{k}^{f}<\frac{1}{2L_{f1}}$.
Based on (27), for constant $0<\alpha_{f}<\frac{1}{2L_{f1}}$ and following the geometric series ($\sum_{k=0}^{\infty}ar^{k}=\frac{a}{1-r},|r|<1$), the following holds:
$\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}$ (28)
$\leq$ $(1-2\alpha_{f}L_{f1})^{k}\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}+\alpha_{f}^{2}L_{f2}\sum_{s=0}^{\infty}(1-2\alpha_{f}L_{f1})^{s}$
$\leq$ $(1-2\alpha_{f}L_{f1})^{k}\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}+\frac{\alpha_{f}L_{f2}}{2L_{f1}}.$
[…]
6.2 Proof of Corollary 1
It can be observed that, by selecting a decaying step size for (28), specifically, $\alpha_{k}^{f}=\frac{\beta_{f}}{2+k}$, where $\beta_{f}=\frac{1}{L_{f1}}$, the following holds:
$\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\leq\frac{\nu_{f}}{2+k},$ (29)
where $\nu_{f}=\max\{\beta_{f}^{2}L_{f2},\quad 2\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\}$. The main idea to prove this is to show $\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\leq\frac{\nu_{f}}{2+k+1}$ using induction. We start from the case when $k=0$, where one has
$\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\leq\frac{\nu_{f}}{2}.$
Next, if $k\geq 1$, one has:
$\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}$ (30)
$\leq$ $(1-2\alpha_{k}^{f}L_{f1})\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}+(\alpha_{k}^{f})^{2}L_{f2}$
$=$ $(1-\frac{2\beta_{f}L_{f1}}{2+k})\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}+\frac{\beta_{f}^{2}L_{f2}}{(2+k)^{2}}$
$\leq$ $\frac{(2+k-1)\nu_{f}}{(2+k)^{2}}+\frac{(1-2\beta_{f}L_{f1})\nu_{f}+\beta_{f}^{2}L_{f2}}{(2+k)^{2}}$
$=$ $\frac{(2+k-1)\nu_{f}}{(2+k)^{2}}+\frac{\beta_{f}^{2}L_{f2}-\nu_{f}}{(2+k)^{2}}\quad(\because\beta_{f}=\frac{1}{L_{f1}})$
$\leq$ $\frac{\nu_{f}}{2+k+1}\quad(\because(x+1)^{2}\geq x(x+2)).$
[…]
6.3 Proof of Lemma 2
To begin, we recall $\hat{J}_{\mu,t^{*}}$ from (17). For simplicity and ease of notation, we denote $\boldsymbol{u}_{t^{*}}^{i}\coloneqq\mu(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}_{i}^{\mu})$ and $\boldsymbol{\hat{x}}_{t^{*}+1}^{i}\coloneqq C_{k}(A_{k}\boldsymbol{g}(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}_{k+1}^{f})+B_{k}\boldsymbol{u}_{t^{*}}^{i})$. By following the definition of $\hat{J}_{\mu,t^{*}}$, for any $\boldsymbol{\theta}_{1}^{\mu},\boldsymbol{\theta}_{2}^{\mu}\in\mathbb{R}^{q}$, the following holds:
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$ (31)
$=$ $\parallel\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{1})\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})+\gamma\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{1})\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{1}$
×$\times\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{2})\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})-\gamma\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{2})$
×$\times\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
$=$ $\parallel\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{1})(\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu}))$
$+(\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{1})-\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{2}))\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})+\gamma\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{1})$
×$\times(\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{1}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu}))$
$+\gamma(\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{1})-\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{2}))\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel.$
Then, by following the norm triangle inequality, subordinance, and submultiplicativity, (31) becomes:
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
$\leq$ $\parallel\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{1})\parallel\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
$+\parallel\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{t^{*}}^{1})-\nabla_{\boldsymbol{u}}c(\boldsymbol{u}_{k}^{2})\parallel\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
$+\gamma\parallel\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{1})\parallel\parallel\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{1}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}$
×$\times\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel+\gamma\parallel\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{1})-\nabla_{\boldsymbol{\hat{x}}}V(\boldsymbol{\hat{x}}_{t^{*}+1}^{2})\parallel$
×$\times\parallel\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}\nabla_{\boldsymbol{\theta}^{\mu}}\boldsymbol{u}_{t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel.$
Finally, given that $\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{1}=\nabla_{\boldsymbol{u}}\boldsymbol{\hat{x}}_{t^{*}+1}^{2}=C_{k}B_{k}$, and by applying the matrix norm bound and Lipschitz continuity, the following result is derived:
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{1}^{\mu})-\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}_{2}^{\mu})\parallel$
$\leq$ $L_{c\mu}L_{\mu\theta\theta}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel+L_{c\mu\mu}L_{\mu\theta}^{2}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel+\gamma L_{vx}L_{\mu\theta\theta}$
×$\times c_{2}c_{3}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel+\gamma L_{vxx}L_{\mu\theta}^{2}(c_{2}c_{3})^{2}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel$
$=$ $(L_{c\mu}L_{\mu\theta\theta}+L_{c\mu\mu}L_{\mu\theta}^{2}+\gamma L_{vx}L_{\mu\theta\theta}c_{2}c_{3}+\gamma L_{vxx}L_{\mu\theta}^{2}(c_{2}c_{3})^{2})$
$\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel$
$\eqqcolon$ $L_{J\theta\theta}\parallel\boldsymbol{\theta}_{1}^{\mu}-\boldsymbol{\theta}_{2}^{\mu}\parallel,$
where $L_{c\mu}$, $L_{c\mu\mu}$ is the Lipschitz constants of the stage cost $c$ and $\nabla c$ with regarding $\boldsymbol{\mu}$, respectively. Similarly, $L_{\mu\theta}$ and $L_{\mu\theta\theta}$ represent the Lipschitz constants of $\boldsymbol{\mu}$ and $\nabla\boldsymbol{\mu}$ with regard to $\boldsymbol{\theta}^{\mu}$, respectively. Moreover, $L_{vx}$ and $L_{vxx}$ denote the Lipschitz constants of $V$ and $\nabla V$ with regarding $\boldsymbol{x}$, respectively. Similarly, one has
$\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}\parallel\leq(L_{c\mu}L_{\mu\theta}+\gamma L_{vx}L_{\mu\theta}c_{2}c_{3})\eqqcolon L_{J\theta}.$
[…]
6.4 Proof of Theorem 1
Recall $\hat{J}_{\mu}=\frac{1}{N}\sum_{t^{*}\in\mathcal{I}_{D}}\hat{J}_{\mu,t^{*}}$ from (15) as the averaged $\hat{J}_{\mu,t^{*}}$ in (17) computed over all $\boldsymbol{x}_{t^{*}}$ in $\mathcal{D}$. The goal of this proof is to analyze the convergence of $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}$ while tuning $\boldsymbol{\theta}^{f}$, $\boldsymbol{\theta}^{J}$, and $\boldsymbol{\theta}^{\mu}$ according to Algorithm 1.
We begin the analysis by employing the Taylor expansion for $\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$ while disregarding higher-order terms, which will lead to the following results:
$\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$ (32)
$=$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})+\langle\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{\mu}-\boldsymbol{\theta}_{k}^{\mu}\rangle$
$+\frac{1}{2}\nabla_{\boldsymbol{\theta}^{\mu}\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel\boldsymbol{\theta}_{k+1}^{\mu}-\boldsymbol{\theta}_{k}^{\mu}\parallel^{2}$
$=$ $-\alpha_{k}^{\mu}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})^{\prime}\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$
$+\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})+\frac{(\alpha_{k}^{\mu})^{2}}{2}\nabla_{\boldsymbol{\theta}^{\mu}\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$
×$\times\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}.$
Building on Lemma 2 and rearranging the results yields the following expression:
$\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$ (33)
$\leq$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$
$+\frac{(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$
$=$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})+\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$
$-\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})+\frac{(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}.$
By applying the Taylor expansion to (33) and considering linear approximators $\boldsymbol{g}$ and $V$, the following expression is obtained:
$\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (34)
$\leq$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k+1}^{\mu})$
$+\langle\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}_{k}^{f}\rangle$
$+\langle\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu}),\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}_{k}^{J}\rangle.$
Here, to simplify the notation, we denote $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})=\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})$, $\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f})=\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$ and $\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{J})=\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f},\boldsymbol{\theta}_{k}^{J},\boldsymbol{\theta}_{k}^{\mu})$ in the rest of this proof. Then taking the summation over the iteration steps $k=0,1,2,\cdots,K-1$ on both sides of (34) and applying Cauchy–Schwarz inequality yields:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (35)
$\leq$ $\sum_{k=0}^{K-1}\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}_{k}^{f}\parallel\parallel\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f})\parallel+\parallel\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}_{k}^{J}\parallel$
×$\times\parallel\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{J})\parallel+\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}_{K}^{f},\boldsymbol{\theta}_{K}^{J},\boldsymbol{\theta}_{K}^{\mu})$
$=$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})+\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})$
$-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}_{K}^{\mu})\!+\!\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}_{K}^{\mu})\!-\!\hat{J}_{\mu}(\boldsymbol{\theta}_{K}^{f},\boldsymbol{\theta}_{K}^{J},\boldsymbol{\theta}_{K}^{\mu})$
$+\sum_{k=0}^{K-1}\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}+\boldsymbol{\theta}^{f*}-\boldsymbol{\theta}_{k}^{f}\parallel\parallel\nabla_{\boldsymbol{\theta}^{f}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{f})\parallel$
$+\parallel\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}^{J*}+\boldsymbol{\theta}^{J*}-\boldsymbol{\theta}_{k}^{J}\parallel\parallel\nabla_{\boldsymbol{\theta}^{J}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{J})\parallel.$
Let $C_{1}=\max\{L_{vx}c_{1}c_{3},1\}$. Following Definition 1, triangle inequality, and Lipschitz continuity, we obtain the following bound:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (36)
$\leq$ $2C_{1}\sum_{k=0}^{K-1}(\parallel\boldsymbol{\theta}_{k+1}^{f}-\boldsymbol{\theta}^{f*}\parallel+\parallel\boldsymbol{\theta}_{k+1}^{J}-\boldsymbol{\theta}^{J*}\parallel)$
$+C_{1}(\parallel\boldsymbol{\theta}_{K}^{J}-\boldsymbol{\theta}^{J*}\parallel+\parallel\boldsymbol{\theta}_{K}^{f}-\boldsymbol{\theta}^{f*}\parallel)$
$+\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*}).$
By selecting the step sizes as $\alpha_{k}^{f}=\frac{\beta_{f}}{2+k}$ from Lemma 1, $\alpha_{J}=(1-\gamma)/4$ from Lemma 3, we derive the following results:
$\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}-\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2}$ (37)
$\leq$ $\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})+C_{1}\Big{(}(1-\nu_{J})^{K/2}$
×$\times\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel+\sqrt{\frac{\nu_{f}}{2+K}}+2\sum_{k=0}^{K-1}(\sqrt{\frac{\nu_{f}}{2+k+1}}$
$+(1-\nu_{J})^{(k+1)/2}\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel)\Big{)}.$
Furthermore, by applying the geometric series formula ($\sum_{k=0}^{\infty}ar^{k}=\frac{a}{1-r},|r|<1$) to (37), and subsequently taking the minimum of the left-hand side of (37), we obtain:
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$ (38)
$\leq$ $\frac{C_{1}}{\sum_{k=0}^{K-1}\alpha_{k}^{\mu}}\Big{(}(\frac{2}{1-\sqrt{1-\nu_{J}}}+(1-\nu_{J})^{K/2})\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel$
$+\sqrt{\nu_{f}}(K^{-1/2}+4K^{1/2})+\frac{1}{C_{1}}(\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})$
$-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*}))\Big{)}+\frac{\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}L_{J\theta}^{2}L_{J\theta\theta}}{2\sum_{k=0}^{K-1}\alpha_{k}^{\mu}}.$
[…]
6.5 Proof of Corollary 2
Choosing the step size $\alpha_{k}^{\mu}=(k+1)^{-1/4}$ for (38), where $\sum_{k=0}^{K-1}\alpha_{k}^{\mu}\approx\int_{k=1}^{K}k^{-1/4}\leq\frac{4K^{3/4}}{3}$ and $\sum_{k=0}^{K-1}(\alpha_{k}^{\mu})^{2}\approx\int_{1}^{K}k^{-1/2}\leq 2\sqrt{K}$, then (38) becomes
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$
$\leq$ $3K^{-1/4}(C_{1}\sqrt{\nu_{f}}+\frac{L_{J\theta}^{2}L_{J\theta\theta}}{4})+\frac{3C_{1}\sqrt{\nu_{f}}K^{-5/4}}{4}$
$+\!\frac{3K^{-3/4}}{4}\Big{(}C_{1}(\frac{2}{1-\sqrt{1-\nu_{J}}}+(1-\nu_{J})^{K/2})\parallel\!\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\!\!\parallel$
$+\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})\Big{)}$
$\eqqcolon$ $L_{a}K^{-1/4}+L_{b}K^{-3/4}+L_{c}K^{-5/4},$
where
$L_{a}$ $=3(C_{1}\sqrt{\nu_{f}}+\frac{L_{J\theta}^{2}L_{J\theta\theta}}{4}),$ (39)
$L_{b}$ $=\frac{3}{4}\Big{(}C_{1}(\frac{2}{1-\sqrt{1-\nu_{J}}}+(1-\nu_{J})^{K/2})\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel$
$\quad+\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})\Big{)},$
$L_{c}$ $=\frac{3C_{1}\sqrt{\nu_{f}}}{4}.$
[…]